OpenStudy (anonymous):

Find f(g(x)) if f(x) = 2x-2 and g(x) = x + 4

OpenStudy (anonymous):

2(x+4)-2=2x+8-2=2(x+3)

OpenStudy (anonymous):

You are awesome! Thanks for the help.

OpenStudy (anonymous):

2x-6

OpenStudy (anonymous):

sorry 2x+6

OpenStudy (anonymous):

take the g(x) equation and place it inside the f(x) equation

OpenStudy (anonymous):

Okay, cool.

OpenStudy (anonymous):

2x+6